# Recommender systems and reinforcement learning for building control and occupant interaction: A text-mining driven review of scientific literature

### v1

@@ 4.2 Well-established applications of recommender systems in building control and occupant interaction
Location recommendation aims to suggest optimal places to users based on their preferences, behaviors, and contextual information [124]. These systems leverage various data sources, including geographic information, user location history, and social networks, to generate personalized location recommendations [125]. For instance, an enhanced recommendation method employs Item-based Collaborative Filtering to recommend new points of interest based on non-spatial attributes similar to previously liked locations [126, 127]. However, traditional collaborative filtering approaches often face the challenge of high sparsity in the user-item matrix, which significantly limits their effectiveness [128]. To address this issue, researchers have proposed methods that integrate various Probability Matrix Factorization models to reduce dependency on specific data types [129]. Additionally, the LCARS methodology proposed in [128] combines offline modeling and online recommendations not only to address the data sparsity issue but also to enhance explainability. In the context of building management, location recommendation systems are also frequently employed for internal space optimization to enhance energy efficiency, comfort, and functionality by optimizing the use of indoor spaces [10]. Given the complexity and diversity of data, methods based on deep learning and reinforcement learning are widely applied in this domain [38, 130]. These algorithms often consider multidimensional factors such as users’ occupancy patterns, environmental conditions, and personal preferences to provide effective solutions for indoor space management and optimization [131].
@@
Personalized control recommendations is another principal application. In the field of built environment, these systems are commonly employed to optimize indoor conditions such as temperature, lighting, and air quality, involving the processing and analysis of time-series data on environmental parameters and user behaviors [10]. Traditional recommendation algorithms, like conventional collaborative filtering, typically lack the capability to handle time-series data and do not possess mechanisms for balancing multiple objectives in complex scenarios [132]. Consequently, personalized control recommendation systems for building energy patterns and indoor environment optimization frequently utilize reinforcement learning, neural networks, and advanced collaborative filtering algorithms to overcome the limitations of conventional recommendation algorithms and address the overload problem [133]. For example, [133] introduced a deep reinforcement learning framework designed to autonomously adjust intelligent lighting systems, enhancing energy efficiency and user satisfaction. Some studies have also replaced collaborative filtering with mathematical models, such as the Manhattan distance, for making recommendations [134]. However, recommender systems face significant challenges in adapting to dynamic environments and balancing multiple objectives [28].
Building energy optimization recommender systems enhance energy efficiency and sustainability by analyzing data from various sources, including occupancy patterns, environmental sensors, and historical energy usage [10]. In the context of global energy challenges, these systems represent a significant future development trajectory. Despite extensive research focused on optimizing building controls, the adoption rates of IoT home appliances and whole-home control systems remain low in residential buildings, leading to limited practical deployment [135]. Consequently, recommender systems that promote behavior change are increasingly viewed as a more effective alternative to conventional smart control systems. According to Figure 11, current methodologies include RL, which dynamically adjusts energy recommendations by learning from interactions with the environment [38]; clustering algorithms, which predict optimal energy settings based on historical data and user behavior [28]; and rule-based models, which offer the advantage of implementing structured decision-making processes that can straightforwardly incorporate expert knowledge and predefined rules into the energy management system [136]. Additionally, optimization algorithms, particularly those integrated with building energy models, are used to address complex multi-objective optimization tasks [137, 138]. The recent focus in this domain includes shifting towards occupant-centered recommendation, integrating personal data to enhance user interactions, designing interaction platforms, and addressing user acceptance and privacy concerns [10].
@@
Community recommendation systems enhance social interaction and community engagement by providing personalized recommendations based on data related to activities, interactions, and preferences [139]. These systems have been widely implemented across various sectors, including community event recommendations, optimization of facility usage, facilitation of resident interactions, and emergency response and safety management [24]. Traditionally, these systems rely on established recommendation algorithms such as collaborative filtering, content-based filtering, and hybrid methods. For instance, [140] proposed cohort comfort models that aims to use occupant similarity to predict an individual’s thermal preference. Furthermore, NLP techniques, such as topic modeling [141], are frequently utilized to enhance the processing of unstructured data, which includes text and conversational data. Additionally, advancements in LLMs, such as GPT (Generative Pre-trained Transformer) APIs, are being leveraged for complex semantic analysis [142]. Ongoing research in this field aims to augment the contextual awareness of these algorithms while addressing challenges related to data privacy and security [143].
@@ 4.3 Emerging applications and opportunities
In Figure 11, the right side illustrates that certain recommendation objectives have notably low correlations with the underlying algorithmic concepts (highlighted with pink border). These weak links suggest that these areas may not be adequately addressed by current algorithmic frameworks, highlighting existing gaps and opportunities for innovation. These objectives include building fault detection and predictive maintenance, product recommendations, and specific indoor environment recommendations such as improving sleep quality and work efficiency.
@@
Automated fault detection and diagnostics (AFDD) has rapidly evolved since the early 1990s [47]. Traditionally, fault detection and diagnosis relied on rule-based algorithms and empirical formulas to generate reactive and preventive maintenance plans [144]. However, with recent advances in AI, the focus has shifted towards ML-based systems, which are now recognized for their ability to identify faults and predict maintenance needs in building systems [145]. These systems exhibit exceptional dynamic adjustment and predictive capabilities, enabling timely maintenance and minimizing downtime [144]. By analyzing sensor data and historical maintenance records, these systems can predict potential failures and recommend maintenance actions before issues escalate, marking a significant departure from the traditional methods of periodic checks or reactive maintenance following failures [146]. For instance, integrating ML algorithms with sensor data from HVAC systems allows for real-time insights into equipment health, thereby facilitating proactive maintenance measures [147].
@@
Product recommendation systems are extensively utilized in e-commerce and retail sectors to enhance the user experience by offering personalized product suggestions through algorithms such as collaborative filtering and content-based filtering [148]. However, the potential application scenarios for these systems in the domains of smart buildings and smart homes remain underexplored and underutilized. Current product recommendations typically consider only user behavioral data or basic information such as purchase history, browsing records, or user reviews [149]. Integrating environmental data and user preferences related to energy savings or thermal comfort can significantly improve the relevance and efficiency of recommendations. For instance, when assisting consumers with purchasing appliances, several factors may influence their choices, including the appliance’s energy consumption, manufacturer, price, and its contribution to improving the indoor environment [10]. Additionally, with the deregulation of the retail electricity market, research has also employed recommender systems to provide electricity customers with personalized recommendations for electricity providers and electricity plans tailored to their individual consumption needs [150]. Furthermore, these recommendations can be extended to various aspects of the built environment, such as building design strategies [151] and energy transition solutions [109].
@@
Specific environment optimization through recommender systems represents a future direction for development. Previous discussions have illustrated that recommender systems are utilized to enhance indoor environments by providing personalized control suggestions, primarily focusing on improving indoor comfort and energy efficiency [10]. Recent research has shifted towards offering optimization recommendations for specific settings based on their unique features, such as enhancing sleep quality and work efficiency in sleeping and working environments [152, 153]. These specialized environments require not only environmental data but also an integration of interdisciplinary domain knowledge and individual physiological data or real-time feedback to furnish more accurate recommendations [154]. With the widespread adoption of smart wearable devices and smart home technologies, the real-time acquisition of such data has become feasible. For instance, the Cozie app [155], based on the Apple Watch and Fitbit, can be utilized to gather micro-ecological momentary assessment data and physiological data. By integrating occupant feedback with other sensor data, personalized adjustment recommendations can be created [156]. However, such highly personalized systems pose challenges in the initial data acquisition, processing, and storage of large volumes of sensitive data [10].
@@ 5 Conclusion
This study employs text mining techniques to perform a comprehensive review of the literature in the context of building control, occupant interaction, and recommendation systems. By extracting and analyzing trends, patterns, and relationships from approximately 27,000 articles sourced from Elsevier journals, findings that are typically elusive to traditional review methods are revealed. The results indicate that recommendation systems have been widely applied to specific content, such as space optimization, location recommendations, and personalized control suggestions. However, the application of recommendation system methodologies to optimize indoor environments and energy efficiency is not as common. This discrepancy is likely due to the fact that these applications often require the integration of interdisciplinary domain knowledge and extensive sensor data. Moreover, while various data sources related to user context, activities, and real-time feedback are frequently utilized, other critical data sources, including those related to physiological and indoor environmental factors, are underutilized despite their importance for optimizing specific environments or objectives. In addition, traditional recommendation algorithms such as collaborative filtering, content-based, and knowledge-based methods are widely used in this area. Lastly, due to the robust data collection capabilities of wearable devices and their user-friendly interfaces, they are commonly utilized as platforms to implement recommendation systems.
@@
In addition, this work addresses several key questions regarding the application of recommender systems in building control and occupant interaction. Specifically, it examines the most commonly utilized algorithms, the most explored applications, and emerging opportunities in this field. Among these, deep reinforcement learning stands out as the predominant algorithm for building-related applications, primarily due to its ability to handle complex inputs and self-update based on occupant feedback. Although certain applications such as location recommendation, personalized control, and energy optimization of buildings are widely adopted, significant opportunities for innovation remain in emerging areas. These include areas such as predictive maintenance of building systems, recommendations for building design, appliances, and electricity plans, as well as specific environment optimization, such as those designed to improve sleep and work productivity based on occupant feedback.
@@
This review also acknowledges some limitations. The methodology may not fully comprehend the nuances and contextual subtleties of academic texts, potentially overlooking the impact of pioneering research. Furthermore, NLP systems might misinterpret terms with multiple meanings or highly specialized terminology without sufficient contextual data. Future directions include integrating transformer-based pertrained models (such as BERT [157]) and LLM, which could improve the comprehension of complex academic texts. Furthermore, fine-tuning these models on specific academic datasets could significantly improve the precision with which key information is extracted, potentially transforming the landscape of automated literature reviews.

### v2

@@ 4.2 Well-established applications of recommender systems in human-building interaction
Location recommendation aims to suggest optimal places to users based on their preferences, behaviors, and contextual information [129]. These systems leverage various data sources, including geographic information, user location history, and social networks, to generate personalized location recommendations [130]. For instance, an enhanced recommendation method employs Item-based Collaborative Filtering to recommend new points of interest based on non-spatial attributes similar to previously liked locations [131, 132]. However, traditional collaborative filtering approaches often face the challenge of high sparsity in the user-item matrix, which significantly limits their effectiveness [133]. To address this issue, researchers have proposed methods that integrate various Probability Matrix Factorization models to reduce dependency on specific data types [134]. Additionally, the LCARS methodology proposed in [133] combines offline modeling and online recommendations not only to address the data sparsity issue but also to enhance explainability. In the context of human-building interaction, location recommendation systems are also frequently employed for internal space optimization to enhance energy efficiency, comfort, and functionality by optimizing the use of indoor spaces [12]. Given the complexity and diversity of data, methods based on deep learning and RL are widely applied in this domain [92, 135]. These algorithms often consider multidimensional factors such as users’ occupancy patterns, environmental conditions, and personal preferences to provide effective solutions for indoor space management and optimization [136].
@@
Personalized control recommendations is another principal application. In the context of human-building interaction, these systems are commonly employed to optimize indoor conditions such as temperature, lighting, and air quality, involving the processing and analysis of time-series data on environmental parameters and user behaviors [12]. Traditional recommendation algorithms, like conventional collaborative filtering, typically lack the capability to handle time-series data and do not possess mechanisms for balancing multiple objectives in complex scenarios [137]. Consequently, personalized control recommendation systems for building energy patterns and indoor environment optimization frequently utilize RL, neural networks, and fusion algorithms to overcome the limitations of conventional recommendation algorithms and address the overload problem [12]. For example, [138] proposed a deep RL-based occupant-centric HVAC control framework. This framework generates personalized HVAC control strategies for different users by analyzing environmental data and user feedback. Additionally, the ReViCEE algorithm has been developed, which analyzes individual and collaborative user preferences from historical data and indoor lighting conditions to generate recommendations aimed at stimulating user engagement and promoting behavioral changes towards energy conservation and sustainability [20]. Recent studies have further explored more personalized recommendations by integrating occupant physiological data with environmental data. For example, [139] has explored calculating fatigue indices from activity data to provide lighting environments that alleviate fatigue. Another study proposed an emotion-oriented recommendation system for occupants, designed to suggest personalized indoor environmental quality settings based on emotional states [19].
Building energy optimization recommender systems enhance energy efficiency and sustainability by providing occupants with tailored energy management suggestions or appliance usage recommendations [12].. In the context of global energy challenges, these systems represent a significant future development trajectory. As shown in Figure 11, RL is the most commonly employed algorithm in this domain. RL demonstrates a unique advantage in its ability to adapt to dynamically changing environments by learning from user behavior and interactions with the surrounding environment [7]. For instance, the RL-based system ”recEnergy” leverages occupant feedback to retrain deep neural networks, thereby improving the effectiveness of energy-saving recommendations and increasing user acceptance rates [92]. In addition, clustering algorithms, which generate energy setting recommendations based on similar cases [140], and knowledge-based algorithms, which produce energy-saving action suggestions by incorporating expert knowledge and predefined rules into energy management systems [12], are also commonly employed due to their practicality. Furthermore, optimization algorithms, especially those integrated with building energy models, are widely used to address complex multi-objective optimization problems [141, 142]. Recent research in this domain has focused on enhancing the temporal awareness of energy recommendation systems. This improvement aims to provide real-time, context-sensitive recommendations that are particularly valuable during emergencies, such as the COVID-19 pandemic [12].
@@
Community recommendation systems are designed to recommend content, services, products, or other resources to community members (such as users or cohort) based on data related to the group’s activities, interactions, and preferences [143]. These systems have been widely implemented across various sectors, including community event recommendations, optimization of facility usage, facilitation of resident interactions, and emergency response and safety management [41]. In the context of occupant comfort, [144] proposed cohort comfort models based on the preferences of similar users to recommend appropriate comfort settings tailored to specific user cohorts. For the algorithm inside, traditionally, these systems rely on established recommendation algorithms such as collaborative filtering and content-based filtering. Additionally, NLP techniques, such as topic modeling [145], are frequently utilized to enhance the processing of unstructured data, which includes text and conversational data. Furthermore, advancements in LLMs, such as GPT (Generative Pre-trained Transformer) APIs, are being leveraged for complex semantic analysis [146]. Ongoing research in this field aims to augment the contextual awareness of these algorithms while addressing challenges related to data privacy and security [147].
@@ 4.3 Emerging applications and opportunities
In Figure 11, the right side illustrates that certain recommendation objectives have notably low correlations with the underlying algorithmic concepts (highlighted with pink border). These weak links suggest that these areas may not be adequately addressed by current algorithmic frameworks, highlighting existing gaps and opportunities for innovation. These objectives include building fault detection and predictive maintenance, product recommendations, and specific environment optimization recommendations.
@@
Automated fault detection and diagnostics (AFDD) has rapidly evolved since the early 1990s [45]. Traditionally, fault detection and diagnosis relied on rule-based algorithms and empirical formulas to generate reactive and preventive maintenance plans [148]. With recent advances in AI, the focus has shifted towards ML-based systems [149]. By analyzing sensor data and historical maintenance records, these systems can predict potential failures and recommend maintenance actions before issues escalate, enabling timely maintenance and minimizing downtime [150]. For instance, integrating ML algorithms with sensor data from HVAC systems allows for real-time insights into equipment health, thereby facilitating proactive maintenance measures [151]. However, the current application of purely data-driven AFDD systems remains limited. This limitation arises from the difficulty and high cost of acquiring fault data. Moreover, in real-world FDD applications, it is possible to encounter unknown system faults that were not accounted for during the training phase [152]. The lack of comprehensive training data prevents these systems from covering all building and system operating conditions effectively. Recently, the advancements in generative AI may address the issue of data scarcity [153]. Additionally, the adoption of Non-Intrusive Load Monitoring (NILM) as a replacement for sub-metering significantly reduces the cost and improves the accuracy of energy consumption data acquisition, particularly at the appliance level [12]. These developments have the potential to substantially lower the cost and complexity of developing anomaly detection systems.
@@
Product recommendation systems are extensively utilized in e-commerce and retail sectors to enhance user experiences by providing personalized product suggestions based on user behavioral data or basic information such as purchase history, browsing records, and user reviews [154, 155]. Recently, studies have highlighted that integrating environmental data with user preferences in product recommendation systems can significantly improve the relevance and effectiveness of recommendations. For instance, [156] designed a recommendation system based on a sensor network that extracts user preferences to suggest indoor comfort products, resulting in a 16% increase in consumer purchases. Additionally, integrating appliance energy consumption data into recommendation systems can guide consumers toward purchasing energy-efficient products, thereby contributing to improved building energy efficiency [12]. For example, a study proposed a recommendation system leveraging NILM techniques to detect appliance usage and provide smart grid users with suggestions for energy-efficient appliances that are most relevant or beneficial to them [157]. Additionally, with the deregulation of the retail electricity market, research has also employed recommender systems to provide electricity customers with personalized recommendations for electricity providers and electricity plans tailored to their individual consumption needs, aiming to help them reduce energy costs [158]. Furthermore, product recommendation systems have also been applied to suggest building design strategies [159] and energy transition solutions [114], offering decision support to occupants.
@@
Specific environment optimization through recommender systems also represents one of the key future direction for development. With the advancement of thermal comfort models, recent research has shifted towards offering optimization recommendations for specific settings based on their unique features, such as enhancing sleep quality and work efficiency in sleeping and working environments [160, 161]. These specialized environments require not only environmental data but also an integration of interdisciplinary domain knowledge and individual physiological data or real-time feedback to furnish more accurate recommendations [162]. With the widespread adoption of smart wearable devices and smart home technologies, the real-time acquisition of such data has become feasible. For instance, the Cozie app [163], based on the Apple Watch and Fitbit, can be utilized to gather micro-ecological momentary assessment data and physiological data. By integrating occupant feedback with other sensor data, personalized adjustment recommendations can be created [164]. However, such highly personalized systems pose challenges in the initial data acquisition, processing, and storage of large volumes of sensitive data [12].
4.4 Implications and future directions of text mining-based literature review
Text mining has become a powerful tool for conducting literature reviews, offering notable advantages that enhance both efficiency and comprehensiveness. Its ability to process large-scale datasets efficiently allows researchers to analyze vast amounts of literature that would be unmanageable with traditional manual methods [45]. This scalability enables the identification of hidden patterns and trends across studies that might otherwise go unnoticed. Additionally, text mining reduces inherent human bias in manual reviews, supporting a more objective synthesis of findings.
Despite its strengths, text mining has limitations. The effectiveness of these techniques heavily depends on the quality and structure of the input data. Noise from poorly managed datasets or insufficient data pre-processing can lead to incomplete or misleading results [71]. Moreover, some text mining methods face challenges in interpretability, making it difficult for researchers to fully understand how specific conclusions are reached, which can undermine the credibility of the findings [165]. Text mining may also overlook nuanced or contextual information, such as subtle implications in methodology or theoretical frameworks, which human reviewers are better equipped to identify[45].
Looking forward, several strategies can enhance the integration of text mining in literature reviews. A promising approach involves combining text mining techniques with traditional review methods, creating a hybrid framework that balances automation with the depth of human expertise. In addition, integrating transformer-based pertrained models, such as BERT [166] and LLMs, can further improve the understanding of complex academic texts [72]. Fine-tuning these models on domain-specific datasets could significantly enhance the precision of extracting key information, revolutionizing automated literature reviews [167]. Finally, developing standardized frameworks for applying text mining in literature reviews can improve reproducibility and comparability across studies, fostering greater consistency across disciplines.
@@ 5 Conclusion
This study employs text mining techniques to perform a comprehensive review of the literature in the context of human-building interaction and occupant context-aware support, and recommendation systems. By extracting and analyzing trends, patterns, and relationships from 27,595 articles sourced from Elsevier journals, findings that are typically elusive to traditional review methods are revealed. Compared to other studies, this study provides the first comprehensive analysis of the applications of recommendation systems in the context of human-building interaction, including energy efficiency, occupant health, and indoor environmental quality. It also quantifies the interrelationships among different components (data sources, algorithm types, system types, platforms, and recommendation objectives) of recommendation systems.
@@
The results indicate that recommendation systems have been widely applied to specific content, such as space optimization, location recommendations, and personalized control suggestions. However, the application of recommendation system methodologies to optimize indoor environments and energy efficiency is not as common. This discrepancy is likely due to the fact that these applications often require the integration of interdisciplinary domain knowledge and extensive sensor data. Moreover, while various data sources related to user context, activities, and real-time feedback are frequently utilized, other critical data sources, including those related to physiological and indoor environmental factors, are underutilized despite their importance for optimizing specific environments or objectives. In addition, traditional recommendation algorithms such as collaborative filtering, content-based, and knowledge-based methods are widely used in this area. Lastly, due to the robust data collection capabilities of wearable devices and their user-friendly interfaces, they are commonly utilized as platforms to implement recommendation systems.
@@
In addition, this work addresses several key questions regarding the application of recommender systems in human-building interaction and occupant context-aware support. Specifically, it examines the most commonly utilized algorithms, the most explored applications, and emerging opportunities in this field. Among these, deep reinforcement learning stands out as the predominant algorithm for building-related applications, primarily due to its ability to handle complex inputs and self-update based on occupant feedback. Although certain applications such as location recommendation, personalized control, and energy optimization of buildings are widely adopted, significant opportunities for innovation remain in emerging areas. These include areas such as predictive maintenance of building systems, recommendations for building design, appliances, and electricity plans, as well as specific environment optimization, such as those designed to improve sleep and work productivity based on occupant feedback.
@@
This review also acknowledges some limitations. The methodology may not fully comprehend the nuances and contextual subtleties of academic texts, potentially overlooking the impact of pioneering research. Furthermore, NLP systems might misinterpret terms with multiple meanings or highly specialized terminology without sufficient contextual data. Future directions include integrating transformer-based pertrained models (such as BERT) and LLM, which could improve the comprehension of complex academic texts. Furthermore, fine-tuning these models on specific academic datasets could significantly improve the precision with which key information is extracted, potentially transforming the landscape of automated literature reviews.